OpenStudy (anonymous):

ugh, I need some help. I'm no good at fractions. Here's the problem. Show all work. Find the product: (-1/3 x^2-1/3)(1/2 x+1/4)

OpenStudy (anonymous):

Okay, so can I get someone to help explain this to me? D:

OpenStudy (anonymous):

i can find the answer but im not sure i can explain the steps

OpenStudy (anonymous):

well, any help would be nice lol

OpenStudy (anonymous):

are you looking to solve for x?

OpenStudy (anonymous):

umm..I think I'm just suppose to find the product..

OpenStudy (anonymous):

or is it a F.O.I.L problem

OpenStudy (anonymous):

lol I'm not sure, wouldn't ask if I knew..

OpenStudy (anonymous):

ok give me just a minute to find the answer

OpenStudy (anonymous):

ok so foil answer wil be -1/6x^3 -1/12x^2-1/6x-1/12

OpenStudy (anonymous):

um..okay.. so how do I get to that answer? D: I need to show my work and would very much like to learn how to work problems like this in the future without any issue..

OpenStudy (anonymous):

ok well F.O.I.L stand s for "first" "outer""inner"&"last" when you have a problem like: (2x-4)(2x+4) you multiply 2x by 2x because it is first of both parenthesis next is the outer which is the first 2x and 4 then multiply inner numbers: -4 and 2x last multiply -4 and 4 together this will be : 4x^2 +6x-6x-16 combine like terms and you will have 4x^2-16 (because the 6s cancel each other out)

OpenStudy (anonymous):

:I so how do I do that with fractions? do I flip the denominators or something? -feels like a complete airhead right now-

OpenStudy (anonymous):

no its the same with fractions. If you use a calculator then do the fractions individually with parenthesis around them.

OpenStudy (anonymous):

I think I got it... this is how it was in my notes.. (-1/3x^2 - 1/3)(1/2x + 1/4) =(-1/3x^2 - 1/3)1/2 x * (-1/2x^2 - 1/3)1/4 =1/2x (-1/2x^2 - 1/3) * 1/4(-1/3x^2 - 1/3) =(1/2x * 1/3x^2)(1/2x * -1/3)(1/4 * -1/3x^2)(1/4 * -1/3) =1/6x^3 - 1/6x -1/12x^2 - 1/12 =1/6x^3 - 1/12x^2 - 1/6x - 1/12